OpenStudy (anonymous):

Integral x^2(e^x^3) - x^2 dx

OpenStudy (anonymous):

Factor out \(x^2\):\[\int x^2\left(e^{x^3}-1\right)~dx\] Now substitute, \(u=x^3\) so \(\dfrac{1}{3}du=x^2~dx\): \[\int \left(e^u-1\right)~du\]

OpenStudy (anonymous):

Thank you (;

OpenStudy (anonymous):

yw